OpenStudy (anonymous):

Find the complete solution in radians of the equation. I really need help on this one. There are pictures provided for the problem and possible answers. Please tell me how you got your answer, and thank you in advance for your help.

OpenStudy (anonymous):

Here is the problem itself.

OpenStudy (anonymous):

And here are the possible answers.

OpenStudy (anonymous):

those are 5 separate problems, not 1 problem and 4 answers

OpenStudy (anonymous):

I'm sorry, but it is a multiple choice question. The first picture I uploaded is the problem and I need to find the complete solution in radians.

OpenStudy (inkyvoyd):

the very first thing to do is convert sin^2 theta into 1-cos^2 theta

OpenStudy (inkyvoyd):

then, factor the equation and solve.

OpenStudy (anonymous):

@inkyvoyd thank you, could you please show me how so I can write it down in my notebook?

OpenStudy (inkyvoyd):

note that sin^2 theta+cos^2 theta=1 therefore sin^2 theta=1-cos^2 theta thus you have cos th (1-cos^2 th)-3cos th /4=0

OpenStudy (inkyvoyd):

then factor the equation for cos th. Tell me what you get.

OpenStudy (anonymous):

@inkyvoyd I'm sorry, but I'm still a little lost. Would it be too much to ask for more help?

OpenStudy (inkyvoyd):

What part are you having trouble understanding?

OpenStudy (anonymous):

The thing is I'm trying to put it into my calculator and I'm having trouble navigating it. Could you show me how to do the equation?

OpenStudy (inkyvoyd):

If you had followed my directions you would not have trouble. The reason you are lost is because you are relying on your calculator. Unless your calculator has CAS capabilities, I believ eyou should follow what I say.

OpenStudy (inkyvoyd):

you cannot solve an algebraic equation by entering it into your calculator. You must first isolate the variable, and even then there is no need for a calculator in this problem, nor will a calculator be helpful.

OpenStudy (inkyvoyd):

What you need to do is factor the equation, as I have mentioned.

OpenStudy (inkyvoyd):

let me know what you get...

OpenStudy (anonymous):

@inkyvoyd I'm sorry, but I don't seem to know how to do this in radians. I tried it out and it came out all strange and not even close to any of the possible choices I have.

OpenStudy (inkyvoyd):

@lalaloopsy8888 , factoring has nothing to do with radians. I'm here to work out this problem step by step with you. If you want to continue, then follow my directions please. Factor the equation as I have given you.

OpenStudy (anonymous):

@inkyvoyd What I got was. cos th/2

OpenStudy (inkyvoyd):

Okay, I'm out of patience amigo. I have waited here for each of your responses, none of which are relevant to what I am talking about. Three strikes and you're out. GL.

OpenStudy (anonymous):

@inkyvoyd Alright, sir, I'm so sorry that I upset you. I'm just really lost and I needed guidance. In no way should my inability to comprehend reflect on my sincerity. Once again, I am so sorry to have upset you.

OpenStudy (inkyvoyd):

Eh, I'm a bit edgy due to lack of sleep. I'll hand the problem over to some one more competent at this time of day. @abb0t @Confusionist @jagatuba and all you useless people in chat, please help this dude

OpenStudy (confusionist):

Inky, you're here to do a job. Do it or leave. >_>

OpenStudy (inkyvoyd):

inky is leaving. Confusionist is in.

OpenStudy (confusionist):

:c You had one job. ONE JOB.

OpenStudy (anonymous):

Thank you for participating. I'm really lost here and I don't even know where to start. Could you please help?

OpenStudy (jagatuba):

Man you need to learn LaTeX or use the equation generator. It's too big of a pain to sift through all those jpegs.

OpenStudy (anonymous):

\[\cos \theta \sin^2 \theta - 3\cos \theta/4=0\]

OpenStudy (anonymous):

This was my question: Find the complete solution in radians of the equation

OpenStudy (perl):

do you want help ?

OpenStudy (perl):

woops left out a solutions

OpenStudy (perl):

I will use t instead of theta. cos t sin^2 t - 3/4 cos t . factor out cos t cos t ( sin^2 t - 3/4) = 0 cos t = 0 or sin^2 t - 3/4 = 0 cos t = 0 or sin^2 t = 3/4 cos t = 0 or sin t = +/- sqrt(3) / 2 t = cos^-1 (0) or t = +/- sin^-1 [ sqrt(3) / 2 ] t = cos^-1(0) or t = sin^-1[ sqrt(3)/2] or t = - sin^-1[sqrt(3)/2] t= { pi/2 + pi*n , pi/3 + 2pi*n , 2pi/3+ 2pi*n, 4pi/3 + 2pi*n, 5pi/3 + 2pi*n}

OpenStudy (anonymous):

@perl Thank you so much! It was so nice of you to help me, I was so confused before. I'll write this procedure down. Thank you!

OpenStudy (anonymous):

@perl Thank you!